OpenStudy (anonymous):

Convert 0.02 g/mL to the unit g/L.

OpenStudy (anonymous):

The "m" in mL stands for milli. Milli denotes the scale of 1/1000. 1 mL = 1/1000 L = 0.001 L To figure out the answer you should use an equation to convert the mL to L. 0.02 g/mL x (1000 mL/ 1 L) = 0.02 x 1000 (g/mL x mL/L) In the above equation, the mL of g/mL and the mL of the mL/L cancel each other out and you are left the correct units. (Write out the equation with the units directly above and below each other and it will become clearer.) To get the correct magnitude (scale of size) of the numbers you are working with, for every 0 in the 1000, bump up the 2 by one decimal place. 0.02 x 1000 = 20

OpenStudy (anonymous):

Thanks